A debt of 40 000 is to amortized with 8 000 being paid at the end of each quarter .The interest rate is 16% compounded quarterly.Construct an amortization schedule

1
Expert's answer
2021-11-30T09:16:20-0500

"PV = PMT \frac{1-(1+i)^{-n}}{i}"


number of payments:

"n=-\frac{ln(1-PVi/PMT)}{ln(1+i)}=-\frac{ln(1-40000\cdot0.04/8000)}{ln(1+0.16/4)}=5.69\approx6"


Interest Paid = i * Remaining Principal

Principal Repaid = Payment Amount – Interest Paid

Remaining Principal = Remaining Principal (from the previous row) – Principal Paid